OpenStudy (woodsymte):

What is the positive root of the equation x^2 + 5x = 150?

Vocaloid (vocaloid):

hint: that equation factors to (x+15)(x-10) so what is the positive root?

OpenStudy (woodsymte):

15?

OpenStudy (woodsymte):

but when I do the equation I come up with 10 as being the positive

Vocaloid (vocaloid):

10 is the positive root

Vocaloid (vocaloid):

remember, you need to set it equal to 0 to find roots (x+15)(x-10) = 0 (x - 10) = 0 x = 10